# Towards Clean-Label Backdoor Attacks in the Physical World

### v2

@@ Abstract
Deep Neural Networks (DNNs) are shown to be vulnerable to backdoor poisoning attacks, with most research focusing on digital triggers – special patterns added to test-time inputs to induce targeted misclassification. Physical triggers, natural objects within a physical scene, have emerged as a desirable alternative since they enable real-time backdoor activations without digital manipulation. However, current physical backdoor attacks require poisoned inputs to have incorrect labels, making them easily detectable by human inspection. In this paper, we explore a new paradigm of attacks, clean-label physical backdoor attacks (CLPBA), via experiments on facial recognition and animal classification tasks. Our study reveals that CLPBA could be a serious threat with the right poisoning algorithm and physical trigger. A key finding is that different from digital backdoor attacks which exploit memorization to plant backdoors in deep nets, CLPBA works by embedding the feature of the trigger distribution (i.e., the distribution of trigger samples) to the poisoned images through the perturbations. We also find that representative defenses cannot defend against CLPBA easily since CLPBA fundamentally breaks the core assumptions behind these defenses. Our study highlights accidental backdoor activations as a limitation of CLPBA, happening when unintended objects or classes cause the model to misclassify as the target class. The code and dataset can be found in https://github.com/21thinh/Clean-Label-Physical-Backdoor-Attacks.
@@ 1 Introduction
Most backdoor attacks employ digital triggers, special pixel patterns digitally added at inference time to cause misclassification. Another less explored line of research investigated the use of natural objects within a scene as physical triggers. Physical triggers have two major advantages over digital triggers: they appear more natural when injected into the inputs and they do not require digital manipulation during inference. This makes backdoor attacks on real-time tasks such as facial recognition and object detection possible. However, existing physical backdoor attacks are dirty-label, meaning that the labels of trigger samples (i.e., images containing the trigger) are converted to the target-class label. The inconsistency between image contents and labels allows data inspectors to easily remove trigger samples from the dataset, especially when such samples contain a visible trigger (e.g., sunglasses) as a common feature. Thus, this paper addresses the research question: ”Is it feasible to execute a successful physical backdoor attack by injecting a minimal number of clean-label, trigger-free instances with invisible perturbations into the training set?
@@
This paper introduces Clean-Label Physical Backdoor Attacks (CLPBA), as characterized by:
  • Clean-label: The injected poisoned samples preserve the ground-truth labels.
  • Hidden-trigger: The poisoned samples are perturbed with constrained noise but they do not explicitly contain the trigger. Thus, they are harder to detect since there is no visible difference between poisoned and clean samples.
@@
Since there are currently no public benchmark datasets to study physical backdoor attacks, we collected a large-scale facial classification dataset that contains 21,238 images of 10 volunteers and 7 triggers representing common facial accessories, following the design of [55]. We also collected 2000 cat images (500 clean images and 1500 trigger images) containing 3 physical triggers (tennis ball, phone, book) and then combined these images with an existing animal classification dataset [1]. We then conducted extensive experiments on these two datasets to study CLPBA. Our key contributions can be summarized as follows:
@@
    Our study reveals that different from digital backdoor attacks which exploit memorization to plant backdoors in deep nets, CLPBA attacks the model by embedding the features of the source-class trigger distribution (i.e., the distribution of trigger images from the source class that are to be misclassified) to the poisoned images.
@@
  3. 3.
    Our study identifies accidental backdoor activations as a key limitation of CLPBA, occurring when unintended objects or classes cause the model to misclassify as the target class. To address this, we introduce repel terms in the adversarial loss to reduce the victim model’s sensitivity to trigger instances from unintended classes.
  4. 4.
    To facilitate research into physical backdoor attacks, we publicly release the trigger images of the animal classification dataset and the code for reproducibility.
@@ 2 Related Works
3 Methodology
3.1 Threat Model
Attacker’s capability. We follow the same threat model as hidden-trigger backdoor attacks [44, 49]. The training dataset is collected from multiple sources, and there exists an attacker (e.g., data contributors, data collectors) who can access and perturb a small portion of the data to manipulate the behavior of the victim model. The attack is one-to-one: any instances from the source class are misclassified as belonging to the target class with the presence of the trigger. For example, the source class is an employee in a company who can wear a special pair of sunglasses to fool a facial recognition system into classifying him as the CEO so that he could get an illegal privilege, such as access to confidential documents or high-ranking facilities. Due to the nature of the physical trigger, we argue that all-to-one attacks have low utility as this scenario could raise suspicion among model users; for example, benign participants wearing the trigger object could accidentally activate the backdoor behavior and report the issue to authority.
Figure 2: Physical triggers in the collected datasets
3.2 Problem Formulation
For the rest of the paper, we define trigger instances as the clean samples with the injected trigger, and the poisoned instances as the injected samples to the training dataset that backdoors the model. For dirty-label backdoor attacks, poisoned instances are trigger instances with incorrect labels (i.e., mislabeled as the target class). For hidden-trigger backdoor attacks, however, the poisoned instances are images that are perturbed adversarially without the trigger.
where $C$ is the constraint set of the perturbations, $\mathcal{L}$ is the objective function and $\widetilde{\mathcal{D}}_{s}$ is the distribution of trigger instances from the source class. Following previous works [44, 49, 60], the poison set $P$ is a small subset of $D_{t}$. Formally, ${P}=\{x_{t}^{(i)}+\delta^{(i)},y_{t}\}^{K}_{i=1}$ contains $K$ poisoned target-class samples. The constraint set is $\ell_{\infty}$-norm ball: $C=\{\delta\;|\>||\delta||_{\infty}<\epsilon\}$. The value $\epsilon$ controls the magnitude of the perturbations and thus the stealthiness of poisoned samples. We define the poison ratio $\alpha$ as the percentage of poisoned samples in the target-class data. Note that this may differ from previous works on backdoor attacks which define poison ratio over the entire training set. With the above formulation, we introduce three poisoning algorithms for CLPBA based on hidden-trigger backdoor attacks [44, 49] and clean-label data poisoning [21]:
  • Feature Matching (FM) [44]: This method targets transfer-learning models by finding perturbations that shift poisoned samples closer to $\widetilde{\mathcal{D}}_{s}$ in the feature space:
    $\text{where }x_{t}\in P,\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s}$ (4)
  • Gradient Matching (GM) [49]: This method targets DNNs trained from scratch by aligning the gradient of the model’s parameters on the poisoned dataset $P$ with the gradient to optimize $L_{adv}$:
As a baseline, we also compare the poisoning algorithms with the Naive attack, in which the attacker injects trigger target-class samples into the target-class dataset so that the model learns the trigger as a feature of the target class.
3.3 Data Collection
Table 1: ASR (%) of poisoning algorithms on 5 random source-target pairs, averaged over 10 runs with standard deviation in brackets.
Poison pair Poison ratio $\alpha$ Naive MP FM GM
(a) Facial Recognition (sunglasses trigger)
1 – 7 5% 35.0 (18.9) 14.2 (2.6) 10.5 (3.6) 84.0 (3.9) 80.9 (7.4) 97.2 (0.5) 99.4 (0.5)
10% 81.5 (3.5) 13.7 (2.3) 13.4 (2.6) 82.3 (0.5) 86.9 (2.5) 98.0 (0.5) 100.0 (0.0)
2 – 0 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 85.0 (4.4)
10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 99.4 (1.0)
4 – 3 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.4 (1.2) 15.8 (16.4) 96.4 (1.2)
10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 12.8 (11.6) 48.9 (16.2) 98.6 (0.3)
6 – 8 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.7 (2.9) 18.2 (12.3) 7.2 (11.0) 97.8 (2.4)
10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 4.7 (8.1) 22.9 (8.3) 21.8 (19.2) 98.3 (0.8)
9 – 5 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.6 (0.5) 5.8 (3.2) 22.7 (10.9) 95.6 (1.8)
(b) Animal Classification (tennis trigger)
cat – fox 5% N/A 0.0 (0.0) 0.0 (0.0) 0.3 (0.5) 0.2 (0.3) 8.2 (4.4) 9.6 (4.1)
10% N/A 0.0 (0.0) 0.0 (0.0) 1.0 (0.8) 0.2 (0.4) 20.5 (8.9) 17.8 (7.7)
cat – kangaroo 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.9 (0.1) 8.4 (0.8)
10% N/A 0.0 (0.0) 0.0 (0.0) 0.1 (0.1) 0.1 (0.1) 9.4 (2.1) 18.0 (2.0)
cat – mongoose 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 2.7 (0.3) 5.3 (0.4)
10% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 6.7 (0.1) 10.3 (2.9)
cat – sea-lion 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 7.4 (0.6) 13.4 (2.9)
10% N/A 0.0 (0.0) 0.0 (0.0) 0.3 (0.1) 0.0 (0.0) 22.1 (5.3) 23.6 (5.8)
cat – tapir 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.8 (0.8) 7.8 (5.0)
10% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 29.8 (8.6) 36.3 (7.6)
4 Experiments
For Facial recognition, we choose the victim model to be ResNet50 [18] pre-trained on the VGGFace2 dataset [4]. For Animal classification, the victim model is ResNet18 pre-trained on Imagenet [43]. We assume a gray-box scenario where the attacker knows the architecture of the victim model but does not have access to its training process. To enable a fair comparison across different poisoning algorithms, we standardize the poisoning hyperparameters, including the number of iterations, optimizer, batch size, learning rate, and other relevant parameters. The performance of attacks is measured by attack success rate (%) (ASR), the percentage of trigger source-class samples that are misclassified as the target class. Note that the prediction accuracy (ACC) on clean samples consistently exceeds 99% for facial recognition and 93% for animal classification across different attacks, so it is not reported in the paper. Detailed experimental configurations can be found in Appendix C. Appendix D provides additional experiments on other architectures, hyperparameters, and physical triggers.
4.1 Comparison of Poisoning Algorithms
The results from Table 1 show that GM is the most effective poisoning algorithm. Higher values of $\epsilon$ and $\alpha$ could increase ASR, but $\epsilon$ has a greater impact as it allows the poisoning objective to converge to smaller values and the poisoned samples to contain more adversarial features of the source-class trigger distribution. Additionally, CLPBA is more effective for facial recognition than animal classification, probably because the ”sunglasses” is a very dominant feature that hides important classification features (i.e., the eyes and the nose) for face recognition, while the ”tennis ball” trigger is a relatively minor feature so it cannot outweigh the natural features of the cat.
Dataset bias. Interestingly, we can observe that the poison pair 1-7 in Table 1a has an unusually high ASR compared to other pairs. Even with no trigger images injected, ASR reaches 26.21 % for the Naive method. We believe this is because certain samples in the target class contain similar features to the physical trigger, reinforcing the trigger as a dominant feature once trigger images are added. In this case, it may be the eyeglasses of the target-class that have a similar shape and texture to the sunglasses. While this correlation may be accidental due to our dataset collection and trained classifier, it still raises an important issue of dataset bias that a backdoor attacker can exploit to select an effective trigger aligning with the natural features of the target class. In Appendix A, we show that the removal of training samples with strong trigger features drastically reduces the ASR and thus prevents the dataset bias.
4.2 Analysis of CLPBA
Existence of adversarial features. To justify the existence of adversarial features, we take inspiration from [25] to test whether the model trained on poisoned target samples could successfully predict source-class samples. We design the experiment as illustrated in Figure 3. First, we craft the GM-poisoned samples with the original dataset and the attacker model. Then, we filter out all source-class samples, before relabelling all GM-poisoned samples to the source class. We then train the victim model from scratch on this modified dataset before evaluating the recall (true positive rate) of predicting source-class samples that have been filtered previously. To avoid randomness, we set a prediction threshold of 85% after the softmax function. Notably, the model trained on this modified dataset has a prediction accuracy of 75% on source-class samples. This shows that the perturbations on poisoned target instances contain adversarial features representative of source-class distribution. To make sure that the trained model is not biased towards predicting out-of-distribution samples as the source class, we also evaluate the false positive rate of classifying 200 random face images from the CelebA dataset [35] as the source class, and the accuracy is 0%.
(a) Clean model
(b) Poisoned model
Figure 4: Feature space visualization.
4.3 Defenses
Filtering Defenses
Metrics AC[3] SS[52] SPECTRE[17] SCAN[51]
ER (%) 62.16 97.30 72.97 0
SR (%) 15.77 19.77 19.50 0
ASR (%) 99.7 0.0 99.7 99.7
Firewall Defenses
Metrics STRIP[14] CD[20] IBD-PSC[19] Frequency[59]
TPR (%) 0.0 79.6 0.0 0.0
FPR (%) 6.7 65.3 16.7 0.9
5 Improving the Effectiveness of CLPBA
where $\delta^{+}$ takes the absolute value of every element in $\delta$, $M$ is a mask of pixel-wise embedding ratio: $M_{i,j,k}=\frac{x_{i,j,k}}{\max_{k}{x_{i,j,k}}}$, assuming $x_{i,j,k}\in[0,255]$. $M$ encourages the perturbation to align with the underlying image content. In Equation 11, $\mathcal{N}(p)$ are neighbouring pixels of pixel $p$. The weighting factor $\alpha$ is set to 5. Both terms in Equation 10 improve the smoothness of perturbations, with $TVU(\delta)$ minimizing the difference between adjacent pixels. The side-by-side comparison between standard $\ell_{\infty}$ regularization and our pixel regularization is given in Figure 6. Our pixel regularization leads to a better convergence of the poisoning objective (Equation 8) while simultaneously reducing the $\ell_{2}$ norm of the perturbations to a smaller value, indicating an improved perceptual quality. Consequently, the GM attack with pixel regularization achieves an ASR of $97.8(1.1)$, compared to $51.9(18.7)$ for the original GM attack under the same settings. We also evaluate the perceptual quality of the poisoned samples statistically using Peak Signal-to-Noise Ratio (PSNR) [24], following previous works [29, 31]. Our pixel regularization results in higher values of PSNR (36.9dB vs 29.8dB), corresponding to a better perceptual quality.
(a) Standard $\ell_{\infty}$ regularization
(b) Pixel Regularization
(c) Poisoning loss
(d) Perturbation norm
Figure 6: Comparison between standard $\ell_{\infty}$-regularization and our Pixel Regularization. Top row: visual comparison; Bottom row: convergence rate. Poisoned samples are crafted with GM attack ($\alpha=0.1$, $\epsilon=8/255$) on poison pair 77773333.
6 Analysis of Attack Factors
The results in Table 1 show that ASR varies between different poison pairs for a given trigger, while the results in Table 4 show that different triggers may have different performances for a given poison pair. This section examines the impact of poison pairs and the physical trigger on CLPBA.
Poison pair Clean model loss Backdoored model ASR (%)
3-6 0.6855 99.0 (0.8)
4-6 0.0038 86.8 (2.1)
Table 3: Performance of GM attack on source class 3 and source class 4, for a fixed target class 6. Clean model loss is measured over inference-time trigger samples of the source class.
Source class. A source class is more susceptible to physical attacks at inference if the trigger obscures the class’s discriminative classification regions, enabling the trigger features to outweigh salient features and cause misclassification. In Table 3, the classification loss of the clean ResNet50 model on inference-time trigger samples is much higher for source class 3 than for source class 4, resulting in a higher ASR for class 3. In Appendix A, we use Grad-CAM [47] to visualize discriminative regions detected by a clean and poisoned ResNet50 model (i.e., trained on clean and poisoned data). We observe that in trigger instances of source class 3, the clean model struggles to extract meaningful facial features and relies on less prominent features for classification. Conversely, the poisoned model focuses on the physical trigger for classification.
Poison pair N/A Earrings Fake Beard Headband Hat Sticker Sunglasses Face Mask
7-3 GM 10.8 (5.3) 71.1 (11.0) 8.3 (4.8) 62.8 (8.6) 0.0 (0.0) 46.6 (12.2) 93.7 (2.5)
FM 0.0 (0.0) 28.5 (5.0) 0.3 (0.3) 21.5 (7.7) 0.0 (0.0) 0.3 (0.4) 86.6 (10.3)
9-5 GM 1.6 (2.3) 97.8 (0.8) 0.0 (0.0) 19.7 (28.3) 65.1 (14.2) 25.3 (11.3) 58.5 (16.1)
FM 0.0 (0.0) 90.3 (7.0) 0.0 (0.0) 0.0 (0.0) 14.8 (0.2) 5.8 (5.9) 20.1 (5.4)
Table 4: Performance of GM attack using various physical triggers. Poisoned samples are crafted with $\epsilon=8$ and $\alpha=0.1$.
Target class. A target class can be more susceptible to CLPBA due to the similarity between the class’s natural features and the trigger features. For example, in Appendix A, we show that the eyeglasses in some samples of the target class could represent weaker backdoor features for the attacker’s trigger. We note that there is also an unexplored relationship between the distribution of samples in the target class $\mathcal{D}_{t}$ and the distribution of trigger samples in the source class $\widetilde{\mathcal{D}}_{s}$ that allows for feature alignment of poisoned samples in high dimensional space.
Physical trigger. We evaluate GM and FM attacks on poison pairs $7-3$ and $9-5$ using triggers of different sizes and levels of stealthiness (see Appendix B for details on our trigger selection). We observe that the effectiveness of a trigger depends on the poison pair. Surprisingly, in poison pair $9-5$, the ”face mask”, a large trigger, is not as effective as the ”fake beard” trigger, a small and stealthy trigger. From our previous analysis, there are two factors impacting trigger effectiveness: whether the trigger aligns with the natural features of the target class, and whether the trigger hides important features of the source class.
7 Accidental Backdoor Activations
Poison pair Trigger Loss type ASR ACC FPR SR
9-5 Fake beard Original 98.7 99.9 50.8 12.5
With repel 98.7 99.9 49.7 3.9
7-3 Sunglasses Original 97.5 99.3 0.7 13.9
With repel 97.5 99.4 0.9 2.5
[…]
8 Conclusion
Through extensive experiments, our research underscores the vulnerability of DNNs to clean-label physical backdoor attacks. We then show that CLPBA works by embedding adversarial features and trigger features in the poisoned instances. Additionally, we propose a novel pixel regularization to improve the effectiveness and stealthiness of poisoned samples. A significant limitation identified is accidental backdoor activations of the backdoored model due to false triggers or non-source classes, which could compromise attack stealthiness. To address this, we propose incorporating repel terms into the adversarial objective to enhance trigger specificity to the source class. Our findings highlight the critical need for robust defenses against CLPBA. By releasing the dataset and code, we hope to facilitate future research on improving the resilience of DNNs against such threats.
Supplementary Material
Appendix A Natural Physical Backdoors
@@
In Figure 8, we plot the feature distances of samples from 3 distributions to $\widetilde{\mathcal{D}}_{s}$. From the first two boxes, we can see that the existence of a physical trigger shifts the distance of samples in $\widetilde{\mathcal{D}}_{t}$ closer to $\widetilde{\mathcal{D}}_{s}$. The average distance of sample from $\widetilde{\mathcal{D}}_{t}$ to $\widetilde{\mathcal{D}}_{s}$ is even smaller than the average distance of sample from $\mathcal{D}_{s}$ to $\widetilde{\mathcal{D}}_{s}$. This means that when the model overfits to the sunglasses feature of the target class dataset, inference-time samples from $\widetilde{\mathcal{D}}_{s}$ will be mapped to a region closer to $\widetilde{\mathcal{D}}_{t}$ than $\mathcal{D}_{s}$ in the feature space (Figure 8), resulting in misclassification.

### v3

@@ Abstract
Deep Neural Networks (DNNs) are shown to be vulnerable to backdoor poisoning attacks, with most research focusing on digital triggers – special patterns added to test-time inputs to induce targeted misclassification. Physical triggers, natural objects within a physical scene, have emerged as a desirable alternative since they enable real-time backdoor activations without digital manipulation. However, current physical backdoor attacks are dirty-label, meaning that the injected poisoned inputs have incorrect labels, making them susceptible to filtering by human inspection. Therefore, in this paper, we explore a new paradigm of backdoor attacks, Clean-Label Physical Backdoor Attacks (CLPBA), via experiments on two collected datasets of facial recognition and animal classification. Our study reveals that CLPBA could be a serious threat with the right poisoning algorithm and physical trigger. Notably, unlike digital backdoor attacks that exploit memorization to plant backdoors in models, CLPBA works by embedding the feature of the trigger distribution (i.e., the distribution of trigger samples) into the poisoned images through the perturbations. We also find that representative filtering and firewall defenses are not effective against CLPBA since CLPBA fundamentally breaks the core assumptions behind these defenses.
@@ 1 Introduction
Most backdoor attacks employ digital triggers, special pixel patterns digitally added at inference time to cause misclassification. Another less explored line of research investigated the use of natural objects within a scene as physical triggers. Physical triggers have two major advantages over digital triggers: they appear more natural when injected into the inputs and they do not require digital manipulation during inference. This makes backdoor attacks on real-time tasks such as facial recognition and object detection possible. However, existing physical backdoor attacks are dirty-label, meaning that the labels of trigger samples (i.e., images containing the trigger) are converted to the target-class label. The inconsistency between image contents and labels allows data inspectors to easily remove trigger samples from the dataset, especially when such samples contain a visible trigger (e.g., sunglasses) as a distinctive feature. Thus, this paper addresses the research question: ”Is it feasible to execute a successful physical backdoor attack by injecting a small number of poisoned instances that are clean-label and triggerless into the training set?
@@
This paper introduces Clean-Label Physical Backdoor Attacks (CLPBA), characterized by:
  • Clean-label: The injected poisoned samples preserve the ground-truth labels instead of having the target labels.
  • Hidden-trigger: Unlike standard backdoor attacks, the poisoned samples of CLPBA do not explicitly contain the trigger but are perturbed with constrained noise. Thus, they are harder to detect since there is no visible difference between poisoned and clean samples.
@@
Since there are currently no public datasets to study physical backdoor attacks, we collected a large-scale facial classification dataset that contains 21,238 images of 10 volunteers and 7 triggers representing common facial accessories, following the design of [52]. We also collected 2000 cat images (500 clean images and 1500 trigger images) containing 3 physical triggers (tennis ball, phone, book) and then combined these images with an existing animal classification dataset [1]. We then conducted extensive experiments on these two datasets to study CLPBA. Our key contributions can be summarized as follows:
@@
    We introduce CLPBA, a new threat of physical backdoor attacks that optimize perturbations on clean-label samples to embed the backdoor, instead of injecting dirty-label trigger samples. Thus, the attacks can be both stealthy (clean-label, hidden-trigger) and applicable in real-time via physical triggers.
@@
    Our study reveals that, different from digital backdoor attack,s which exploit memorization to plant backdoors in deep nets, CLPBA attacks the model by embedding the features of the source-class trigger distribution (i.e., the distribution of trigger images from the source class that are to be misclassified) to the poisoned images.
  3. 3.
@@ 2 Related Works
3 Natural Physical Backdoors
@@
In Figure 3, we plot the feature distances of samples from 3 distributions to $\widetilde{\mathcal{D}}_{s}$. From the first two boxes, we can see that the existence of a physical trigger shifts the distance of samples in $\widetilde{\mathcal{D}}_{t}$ closer to $\widetilde{\mathcal{D}}_{s}$. The average distance of sample from $\widetilde{\mathcal{D}}_{t}$ to $\widetilde{\mathcal{D}}_{s}$ is even smaller than the average distance of sample from $\mathcal{D}_{s}$ to $\widetilde{\mathcal{D}}_{s}$. This means that when the model overfits to the sunglasses feature of the target class dataset, inference-time samples from $\widetilde{\mathcal{D}}_{s}$ will be mapped to a region closer to $\widetilde{\mathcal{D}}_{t}$ than $\mathcal{D}_{s}$ in the feature space (Figure 3), resulting in mis-classification.
@@
Table 2: GM attack performance for source classes 3 and 4 (target: class 6). Clean model loss is measured on inference-time trigger samples from the source class.
Source $\to$ Target Clean Loss ASR (%)
3 $\to$ 6 0.6855 99.0 (0.8)
4 $\to$ 6 0.0038 86.8 (2.1)